Find the values of x and y that solve the following system of equations
8x+3y=-13
4x-7y=-15

Answer by mananth(16946)

8x+3y=-13
4x-7y=-15 Multiply this equation by 2 and subtract the abobe equation
2(4x-7y)-(8x+3y)=2*-15-(-13)
8x-14y-8x-3y=-17
-17y=-17
y=1
plug the value of y in the equation
x=-2